OpenStudy (anonymous):

please help!!! please explain how to get domain and range for y=-2(x-2)^2+1 show complete solution please

OpenStudy (sbuck98):

Tigeralgebra.com gives you answers.! Medal me please!

OpenStudy (anonymous):

i need more help then that lol

OpenStudy (sbuck98):

no like it seriously gives you the exact answer worked out

OpenStudy (sbuck98):

it shows you all the work to it.

OpenStudy (nincompoop):

what does it mean when you're looking for domain and what does it mean to look for range?

OpenStudy (anonymous):

yes

OpenStudy (marcelie):

domain refers to x axis only and range lowest to top on y axis

OpenStudy (nincompoop):

usually for a continuous function, the domain is all Real Numbers \(\mathbb{R} \) numbers Think of the domain as possible values for x, and the range would be the output for y for each x.

OpenStudy (anonymous):

i need to explain how to get the domain and range for this function

OpenStudy (nincompoop):

First, identify if your equation or function is continuous. Do you know how to do that?

OpenStudy (anonymous):

Domain: (-∞, ∞) Range: [-100, ∞) but i dont know how to explain on how to get it

OpenStudy (nincompoop):

how'd you get the -100 for the range ?

OpenStudy (anonymous):

do we not plug in zero?

OpenStudy (marcelie):

range : ( - infinity , 1)

OpenStudy (nincompoop):

that is if your x is zero, but what if it isn't zero?

OpenStudy (anonymous):

ohhhh so sorry, i wrote the wrong equation:(

OpenStudy (nincompoop):

you indicated in your domain a negative infinity, so how are you getting -100 ? how did you evaluate it?

OpenStudy (nincompoop):

let us try a simpler concept before we delve into functions or equations, do you mind?

OpenStudy (anonymous):

its y=lxl-100

OpenStudy (nincompoop):

LAUGHING HYSTERICALLY INSIDE MY HEAD

OpenStudy (nincompoop):

so that is your function? \(f(x) = |x| - 100 \)

OpenStudy (anonymous):

yes

OpenStudy (nincompoop):

so how about we evaluate the function

OpenStudy (anonymous):

okay, how do we d this?

OpenStudy (nincompoop):

before we proceed, have you encountered absolute values before?

OpenStudy (nincompoop):

hello? echoooooo hellooooooo o o o o o

OpenStudy (nincompoop):

omg here's math male he's going to invade this thread

OpenStudy (mathmale):

Ladies and gentlemen! Seems as though we are talking about two different functions here! Unless I'm sorely mistaken, we are to find the domain and range for the quadratic function y=-2(x-2)^2 + 1. Is that correct, or not?

OpenStudy (nincompoop):

it is not correct

OpenStudy (anonymous):

sorry, ive not got much time to explain this and having difficult time as i cannot explain in eng

OpenStudy (nincompoop):

she said she put the wrong function initially and then gave the proper one

OpenStudy (nincompoop):

okay, I am trying to figure out if you have a good idea how to interpret absolute values so I do not start with something you haven't encountered

OpenStudy (mathmale):

Oh, OK. So you want to focus on the problem involving absolute value. All right. I'd suggest that you, @Sammii09, graph y = |x| first. Doing so will help you to understand domain and range in connection with the abs. val. function.

OpenStudy (anonymous):



OpenStudy (nincompoop):

that is a good way to show the general graph of absolute value

OpenStudy (nincompoop):

there should be a symmetry of values from the -x and +x and also the same value of +y

OpenStudy (nincompoop):

now that is graphically, how about we do it analytically

OpenStudy (anonymous):

i know how to graph, i just need to help me explain how to get domain and range please

OpenStudy (mathmale):